tims age is 6 more than twice kims age, the differnce between their ages is 14. how old are tim and kim?
Answer by ReadingBoosters(3246)

x = tim
y = kim
x = 6 + 2(y)
x - y = 14
Substitute the first equation into the second
6 + 2y - y = 14
Combine like terms
y + 6 - 6 = 14 - 6
y = 8
Plu y into the first equation
x = 6 + 2y = 6 + 2(8) = 6 + 16 = 22
Tim is 22
Kim is 8